Burns Company's 2019 and 2018 balance sheet presented the following data for equipment. Equipment 12/31/2019 $325,000 12/31/2018 $262,00. Accumulated depreciation 21/31/2019 $133,900 12/31/2018 $108,000. Book value 12/31/2019 $191,100 12/31/2018 $154,000. During 2019, equipment costing $41,000 with accumulated depreciation of $36,700 was sold for cash, producing a $3200 gain. 1. Calculate the amount of depreciation expense for 2019. 2. Calculate the amount of cash spent for equipment during 2019. 3. Calculate the amount that should be included as a cash inflow from the disposal of equipment.
Answer #1

1) Amount of depreciation expense = (133900-108000)+36700 = $62600

2) Cash spending for equipment = 262000-41000-325000 = $104000

3) Cash inflow from the disposal of equipment = 41000-36700+3200 = 7500